6. If Z is N(0, 1), find values of c such that: (a) Pr(Z> c)=.025 96 (> Iz1)-d (q) (c) Pr(Z> c).05 (d) Pr(Z < c)= 9
Answer #1

The value of Z corresponding to the given probabilities can be obtained from the standard normal distribution table.

a) P(Z \small \geq c) = 0.025

P(Z < c) = 1 - 0.025

P(Z < c) = 0.975

c = 1.96

b) P(|Z| \small \leq c) = 0.95

P(-c < Z < c) = 0.95

P(Z < c) = 0.95 + (1 - 0.95)/2

P(Z < c) = 0.975

c = 1.96

c) P(Z > c) = 0.05

P(Z < c) = 1 - 0.05

P(Z < c) = 0.95

c = 1.645

d) P(|Z| \small \leq c) = 0.9

P(-c < Z < c) = 0.9

P(Z < c) = 0.9 + (1-0.9)/2

P(Z < c) = 0.95

c = 1.645